OpenStudy (anonymous):

how would you find the equation of the straight line that is perpendicular to the line 2x+3y-8= 0 at the point (1,2).?

OpenStudy (hang254):

2(1)+3(2)-8=0 2+6-8=0 8-8=0 0

OpenStudy (dumbcow):

use the slope-intercept form of a line: y = mx+b and perpendicular lines have opposite reciprocal slopes

OpenStudy (dumbcow):

start by finding the slope of the given line

OpenStudy (anonymous):

the slope of original formula is 2/3 the opposite reciprocal would be -3/2 right? so how would i find b?

OpenStudy (dumbcow):

you find b or "y-intercept" by plugging the given point (1,2) in for x and y also i think you are off by a negative....slope should be -2/3

OpenStudy (dumbcow):

y = mx+b if m=3/2 --> 2 = (3/2)(1) + b solve for b

OpenStudy (anonymous):

thank you:)